# Quantum mixing on large Schreier graphs

### v1

@@ 5.2 Main assumptions
When $\Gamma$ has property RD with constant $C_{1}$, the analog of assumption (1.15) is that for some $C^{\prime}_{1}>2C_{1}+1$,
@@
Our definitions of quantum ergodicity (1.8) and (small-scale) quantum (weak) mixing (1.11) are then considered with this definition of the average $\langle K_{N}\rangle$. If $K_{N}$ is a diagonal matrix, its average $\langle K_{N}\rangle$ is a multiple of the identity $\text{Id}_{r}\otimes\text{Id}_{N}$.
@@ 5.3 Results on quantum mixing
the first line is subadditivity of norms, the second line is property RD for $\Gamma$ and the last line is Cauchy-Schwarz. Similarly, if $\rho_{N}$ converges strongly toward $\lambda$ as defined in Definition 1.6 then for any $p\in M_{r}(\mathbb{C})[\Gamma]$, we have
@@
where $f_{j}=f_{j,\eta}$ is a well-chosen polynomial such that $f_{j}\geq 0$ and $f_{j}(\lambda)\geq 1$ on $J^{\eta}_{E_{j}}$. Exactly, as in Section 4.1, we choose $f_{j}=4q_{z_{j},n}$ with $z_{j}=E_{j}+i\eta$ and $\eta\epsilon\geq c/n$ with $q_{z,n}$ and $c$ as in Lemma 2.3 and $\epsilon$ given by (4.3). In all estimates, benign polynomial factors in $r$ will appear in the various constants $C>0$ of Section 4.
@@
We pick $\eta_{N},\epsilon_{N}$, $n_{N},\ell_{N}$ and $\mathcal{E}_{\eta_{N}}$ as in the proof of Theorem 1.5. As explained above (4.15), it is enough to prove that for each $i\in\left[r\right]^{4}$, there exists a vector subspace $H_{N,i}\subset\mathbb{C}^{N}$ of dimension $\mathrm{dim}(H_{N,i})=o(N)$ such that for any $f\in H_{N,i}^{\perp}$ with $\|f\|_{2}\leq 1$, we have
$\lim_{N\to\infty}\max_{E_{1},E_{2}\in\mathcal{E}_{\eta_{N}}}\frac{\|Q_{N,i}f\|_{2}}{\eta_{N}}=0,$ (5.15)
where we recall that $Q_{N,i}$ depends implicitly on $(t,E_{1},E_{2},\eta_{N})$. Indeed, if (5.15) holds true, we set $\bar{H}_{N}=\sum_{i\in[r]^{4}}H_{N,i}\subset\mathbb{C}^{N}$ and define $H_{N}=\{M\in M_{r}(\mathbb{C})\otimes M_{N}(\mathbb{C}):\hbox{for all }(i_{1},i_{2})\in\left[r\right]^{2}:M_{i_{1}i_{2}}\in\bar{H}_{N}\}$ and Theorem 5.3(ii) follows for this choice of $H_{N}$.
@@
In particular, for each $E_{1},E_{2}\in I$, the vector space $H_{N,i,E_{1},E_{2}}$ spanned by eigenvectors of $|Q_{N,i}|=|\rho_{N}(q_{i,t,E_{1},E_{2},\eta_{N}})|$ associated with an eigenvalue larger that $2\|Q_{i}\|_{\rm op}$ has dimension at most $|{\rm Bad}_{S}(n_{N}\ell_{N})|+N2^{-2\ell_{N}}$. Let $H_{N,i}=\sum_{E_{1},E_{2}\in\mathcal{E}_{\eta_{N}}}H_{N,i,E_{1},E_{2}}$. This vector space has dimension at most $o(N)$ by our choice of parameters as explained in the proof of Theorem 1.5. Finally, if $f\in H_{N,i}^{\perp}$ then by construction, for all $E_{1},E_{2}\in\mathcal{E}_{\eta_{N}}$, we have
@@ 6.1.2 Properties of the resolvent
By construction, for any $x\in\Gamma\setminus\{e\}$, there is a unique $k\in\mathbb{N}$ and a unique sequence $(x_{j})_{1\leq j\leq k}$, $x_{j}\in\Gamma_{i_{j}}\setminus\{e_{i_{j}}\}$, such that $i_{j}\neq i_{j-1}$ and $x=x_{k}\ldots x_{1}$. The group element $x$ is said to have length $k$ in the word metric, and $x=x_{k}\ldots x_{1}$ is its reduced expression. The following lemma is a generalization of [77].
@@
Let $z\in\mathbb{H}$ and $x\in\Gamma$ with reduced expression $x=x_{k}\ldots x_{1}$. Let also $y=v_{\ell}\ldots v_{1}x$, where $\ell\geq 1$ and for any $j\in[\ell]$, $v_{j}\in\Gamma_{i_{j}}\setminus\{o_{i_{j}}\}$ for some $i_{j}\in[m]$. We also assume that $y$ is reduced, meaning that $i_{j}\neq i_{j-1}$ and $x_{k}\notin\Gamma_{i_{1}}$. Let $y_{0}=x$ and $y_{j}=v_{j}\ldots v_{1}x$ for $j\in[\ell]$. Then
@@
In the above statement, $y_{j}$ is not necessarily a neighbor of $y_{j-1}$. This depends whether $p(y_{j-1},y_{j})\neq 0$ or not. From now on we denote by $V_{i}$ the set of $x\in\Gamma$ whose reduced expression $x=x_{k}\ldots x_{1}$ satisfies $x_{k}=i$. Lemma 6.4 applies only to $x,y$ in different $V_{i}$. For $x,y$ in the same $\Gamma_{i}$, we have the following expansion instead. Here and below we see $\Gamma_{i}$ as a subset of $\Gamma$, and elements of $\Gamma_{i}$ have word length 0 or $1$ in $\Gamma$.
@@
This relation is derived in [77, Proposition 3] in specific cases. Note that $p(y,y)=0$ if there are no self-loops, i.e. $p_{e}=0$, and that we do not assume $p(y,x)\neq 0$.
@@
Let $n_{i}=|\Gamma_{i}|$. To each $\Gamma_{i}$ we associate the $\nu_{i}=n_{i}(n_{i}-1)$ functions $z\mapsto\zeta_{x}^{z}(y)$ for $(x,y)\in\Gamma_{i}\times\Gamma_{i}$ with $x\neq y$. [^4] Actually we have $\zeta_{x}^{z}(y)=\zeta_{e}^{z}(yx^{-1})$, so there are only $n_{i}$ distinct functions, but for our argument, there is no need to use this reduction. Ordering them arbitrarily, we denote them by $\zeta_{i,j,k}$, where $i\in[m]$, $j\in[n_{i}]$ and $k\in[n_{i}-1]$. Equation (6.9) thus gives the following finite system of quadratic equations, of which $\{h_{i,j,k}=\zeta_{i,j,k}^{z}\}$ (with $i\in[m]$, $j\in[n_{i}]$, and $k\in[n_{i}-1]$) is a solution: for any $i\in[m]$, $j\in[n_{i}]$, and $k\in[n_{i}-1]$,
@@
The statement in Proposition 6.6 appeared in a number of works, in various degrees of generality and justification [8, 43, 77], sometimes by simply exhibiting a system of algebraic equations that $\zeta_{x}^{z}(y)$ satisfies. This is not enough in general, for example $h_{1}+zh_{2}=0$ and $2h_{1}+2zh_{2}=0$ has a solution $h_{1}(z)=ze^{z}$ and $h_{2}(z)=-e^{z}$ which is not algebraic.
@@
The spectrum $\sigma(\lambda(p))$ has a non-trivial absolutely continuous part. There exists a discrete set $D_{0}\subset\mathbb{R}$ such that for any $E\in\mathbb{R}\setminus D_{0}$, $i\in[m]$ and $x,y\in\Gamma_{i}$, the function $\eta\mapsto\zeta_{x}^{E+i\eta}(y)$ has a finite limit as $\eta\rightarrow 0$, and the functions $E\mapsto\zeta_{x}^{E+i0}(y)$ are analytic on $\mathbb{R}\setminus D_{0}$. Furthermore, the maps $E\in\sigma(\lambda(p))\setminus D_{0}\mapsto R^{E+i0}(e,e)$ and $E\in\sigma(\lambda(p))\setminus D_{0}\mapsto\zeta_{x}^{E+i0}(y)$ have finitely many zeroes.
@@
Next, we know from Proposition 6.6 and the Newton-Puiseux theorem [69, Theorem 3.5.2] that there exists a discrete set $D_{0}\subset\mathbb{R}$ such that for any $i\in[m]$ and $x,y\in\Gamma_{i}$, the map $z\mapsto\zeta_{x}^{z}(y)$ is analytic near any $z_{0}\notin D_{0}$ (the function has moreover an expansion as a Puiseux series near $z_{0}\in D_{0}$, and the set $D_{0}$ is necessarily real because we already know that $\zeta_{j,\ell}^{z}$ is analytic on $\mathbb{C}\setminus\mathbb{R}$). The function $\zeta_{x}^{z}(y)$ is also nonzero on $\mathbb{C}\setminus\mathbb{R}$ by definition. It follows from analyticity that the zeroes of $\zeta_{x}^{z}(y)$ are in some discrete set $D_{0}\cup D_{1}$, thus inside the compact set $\sigma(\lambda(p))$ there are only finitely many of them.
We next prove the statement for $R^{z}$. As the function $E\mapsto\sum_{y}p(x,y)\zeta_{x}^{E+i0}(y)-E$ is analytic on $\mathbb{R}\setminus D_{0}$, its zeroes form a discrete set $F_{1}$, and its intersection with $\sigma(\lambda(p))$ which is compact is finite. By (6.4), we deduce that $E\mapsto R^{E+i0}(e,e)$ is analytic on $\mathbb{R}\setminus(D_{0}\cup F_{1})$. Note that $R^{E+i0}(e,e)$ is necessarily nonzero on $\mathbb{R}\setminus D_{0}$ by (6.4). ∎
@@
The spectrum $\sigma(\lambda(p))$ of a free product as above consists of a union of intervals $J_{j}$ of absolutely continuous spectrum, plus possibly finitely many eigenvalues, and (1.5) holds in any closed interval $I_{0}$ contained in the interior of some $J_{j}$ and does not contain any eigenvalue of $\lambda(p)$.
Arguing as in [7] one could prove that the absolutely continuous part is a finite union of intervals, but we shall not pursue this here.
@@
Following the notation introduced in the proof of Proposition 6.7, let $F=D_{0}\cup D_{1}\cup F_{1}$. Since $E\mapsto R^{E+i0}(e,e)$ is analytic on $\mathbb{R}\setminus F$, the set $M^{ac}=\{E\in\mathbb{R}\setminus F:0<\Im R^{E+i0}(e,e)<\infty\}$ is non-empty, and open. Thus, $M^{ac}$ is a union of open intervals, and the absolutely continuous part $\overline{M^{ac}}$ of the spectrum is a union of intervals $J_{j}$. The eigenvalues of $\lambda(p)$ are poles of $E\mapsto\Im R^{E+i0}(e,e)$, which occur only if $E\in F$, and since $\lambda(p)$ is bounded, $\sigma(\lambda(p))\cap F$ is a finite set. By definition and continuity of $E\mapsto\Im R^{E+i0}(e,e)$, we know that (1.5) holds true on each closed interval $I_{0}$ contained in the interior of some $J_{j}$. Such $I_{0}$ is compact as $\sigma(\lambda(p))$ is bounded. ∎
@@
In this section, we check that (1.15) holds. We first prove an elementary lemma which shows that in order to prove the $4$-th moment bound (1.15) in regions where (1.5) holds, it is sufficient to check the convergence of a quantity involving a supremum, instead of a sum.
@@
$\sum_{g\in\Gamma}\eta^{2}(\Im R_{p}^{\lambda_{0}+i\eta})(e,g)^{4}|g|^{C_{1}^{\prime}}\leq\Bigl(\sup_{g\in\Gamma}\eta|(\Im R_{p}^{\lambda_{0}+i\eta})(e,g)|^{2}|g|^{C_{1}^{\prime}}\Bigr)\sum_{g\in\Gamma}\eta(\Im R^{\lambda_{0}+i\eta}_{p})(e,g)^{2}.$
@@
We turn to the proof of (1.15). Let us show that there exists $\delta>0$ such that for any $E\in I$, any $\eta$ small enough, any distinct $i_{1},i_{2}\in[m]$, and any $j_{1},k_{1}\in[n_{i_{1}}],j_{2},k_{2}\in[n_{i_{2}}]$ with $j_{\ell}\neq k_{\ell}$ for $\ell=1,2$, there holds
@@
$\omega_{\underline{\ell}}^{z}=\omega_{\ell_{1}}^{z}\ldots\omega_{\ell_{r}}^{z}.$
Finally for $s=0,1,2$, we pick $g_{s}\in\Gamma_{i_{s}}$ such that $\zeta^{z}_{i_{s},j_{s},k_{s}}=\zeta^{z}_{e}(g_{s})$ (for this, we first pick $x_{s},y_{s}$ such that $\zeta^{z}_{i_{s},j_{s},k_{s}}=\zeta^{z}_{x_{s}}(y_{s})$ and then let $g_{s}=y_{s}x_{s}^{-1}$). Since $\ell_{i}\in\{0,1\}$, it follows that $g_{\ell_{r}}g_{2}g_{\ell_{r-1}}g_{2}\ldots g_{\ell_{1}}g_{2}\in\Gamma$, where $g_{2}$ appears every other time. Therefore we deduce by Lemma 6.4 and the Ward identity (2.16)
@@
Write any $y\in\Gamma$ as $y=g_{i_{r},j_{r}}\ldots g_{i_{1},j_{1}}$ where $g_{i_{\ell},j_{\ell}}\in\Gamma_{i_{\ell}}$ and $i_{\ell}\neq i_{\ell+1}$ for any $\ell\in[r-1]$. Set $L(y)=k$. We have $L(y)\geq|y|/M$ where $M=\sup_{i}|\Gamma_{i}|$ and recall that $|y|$ denotes the word metric of $y\in\Gamma$. Combining (6.12) with Lemma 6.4, we obtain

### v2

@@ 5.2 Main assumptions
When $\Gamma$ has the RD property with constant $C_{1}$, the analog of assumption (1.15) is that for some $C^{\prime}_{1}>2C_{1}+1$,
@@
Our definitions of quantum ergodicity (1.8) and (small scale) quantum (weak) mixing (1.11) are then considered with this definition of the average $\langle K_{N}\rangle$. If $K_{N}$ is a diagonal matrix, its average $\langle K_{N}\rangle$ is a multiple of the identity $\text{Id}_{r}\otimes\text{Id}_{N}$.
@@ 5.3 Results on quantum mixing
the first line is subadditivity of norms, the second line is property RD for $\Gamma$ and the last line is Cauchy-Schwarz. Similarly, if $\rho_{N}$ converges strongly toward $\lambda$ as in Definition 1.6, then for any $p\in M_{r}(\mathbb{C})[\Gamma]$, we have
@@
where $f_{j}=f_{j,\eta}$ is a well-chosen polynomial such that $f_{j}\geq 0$ and $f_{j}(\lambda)\geq 1$ on $J^{\eta}_{E_{j}}$. Exactly, as in Section 4.1, we choose $f_{j}=4s_{z_{j},n}$ with $z_{j}=E_{j}+i\eta$ and $\eta\epsilon\geq c/n$ with $s_{z,n}$ and $c$ as in Lemma 2.3 and $\epsilon$ given by (4.3). In all estimates, benign polynomial factors in $r$ will appear in the various constants $C>0$ of Section 4.
@@
We pick $\eta_{m},\epsilon_{m}$, $n_{N},\ell_{N},f_{N}$ and $\mathcal{E}_{\eta_{m}}$ as in the proof of Theorem 1.5. As explained above (4.15), it is enough to prove that for each $i\in\left[r\right]^{4}$, there exists a vector subspace $H_{N,i}\subset\mathbb{C}^{N}$ of dimension $\mathrm{dim}(H_{N,i})=o(N)$ such that for any $f\in H_{N,i}^{\perp}$ with $\|f\|_{2}\leq 1$, we have
$\lim_{m\rightarrow+\infty}\lim_{N\to\infty}\max_{E_{1},E_{2}\in\mathcal{E}_{\eta_{m}}}\frac{\|Q_{N,i}f\|_{2}}{N\eta_{m}}=0,$ (5.15)
where we recall that $Q_{N,i}$ depends implicitly on $(t,E_{1},E_{2},\eta_{m})$. Indeed, if (5.15) holds true, we set $\bar{H}_{N}=\sum_{i\in[r]^{4}}H_{N,i}\subset\mathbb{C}^{N}$ and define $H_{N}=\{M\in M_{r}(\mathbb{C})\otimes M_{N}(\mathbb{C}):\hbox{for all }(i_{1},i_{2})\in\left[r\right]^{2}:M_{i_{1}i_{2}}\in\bar{H}_{N}\}$ and Theorem 5.3(ii) follows for this choice of $H_{N}$.
@@
In particular, for each $E_{1},E_{2}\in I$, the vector space $H_{N,i,E_{1},E_{2}}^{(m)}$ spanned by eigenvectors of $|Q_{N,i}|=|\rho_{N}(q_{i,t,E_{1},E_{2},\eta_{m}})|$ associated with an eigenvalue larger that $2\|Q_{i}\|_{\rm op}$ has dimension at most $|{\rm Bad}_{S}(n_{N}\ell_{N})|+N2^{-2\ell_{N}}$. Let $H_{N,i}^{(m)}=\sum_{E_{1},E_{2}\in\mathcal{E}_{\eta_{m}}}H_{N,i,E_{1},E_{2}}^{(m)}$. This vector space has dimension at most $C^{2}4^{m}\frac{N}{f_{N}}$ by our choice of parameters as explained in the proof of Theorem 1.5. For $N\in\mathbb{N}$, we finally form the sum of vector spaces $H_{N,i}=\sum_{1\leq m\leq\alpha_{N,i}}H_{N,i}^{(m)}$ where $\alpha_{N,i}$ is the largest integer such that $\sum_{1\leq m\leq\alpha_{N,i}}\dim(H_{N,i}^{(m)})\leq N/\sqrt{f_{N}}$. Finally, if $f\in H_{N,i}^{\perp}$ then by construction, for all $E_{1},E_{2}\in\mathcal{E}_{\eta_{m}}$, we have
@@ 6.1.2 Properties of the resolvent
By construction of the free product, for any $x\in\Gamma\setminus\{e\}$, there is a unique $k\in\mathbb{N}$ and a unique sequence $(x_{j})_{1\leq j\leq k}$, $x_{j}\in\Gamma_{i_{j}}\setminus\{e_{i_{j}}\}$, such that $i_{j}\neq i_{j-1}$ and $x=x_{k}\ldots x_{1}$. The group element $x$ is said to have length $k$ in the word metric, and $x=x_{k}\ldots x_{1}$ is its reduced expression. The following lemma is a generalization of [77, Lemma 4].
@@
Let $z\in\mathbb{H}$ and $x\in\Gamma$ with reduced expression $x=x_{k}\ldots x_{1}$. Let also $y=v_{\ell}\ldots v_{1}x$, where $\ell\geq 1$ and for $j\in[\ell]$, $v_{j}\in\Gamma_{i_{j}}\setminus\{e\}$ for some $i_{j}\in[m]$. We also assume that $y$ is reduced, meaning that $i_{j}\neq i_{j-1}$ and $x_{k}\notin\Gamma_{i_{1}}$. Let $y_{0}=x$ and $y_{j}=v_{j}\ldots v_{1}x$ for $j\in[\ell]$. Then
@@
In the above statement, $y_{j}$ is not necessarily a neighbor of $y_{j-1}$. This depends whether $p(y_{j-1},y_{j})\neq 0$ or not. From now on we denote by $V_{i}$ the set of $x\in\Gamma$ whose reduced expression $x=x_{k}\ldots x_{1}$ satisfies $x_{k}\in\Gamma_{i}$. Lemma 6.4 applies only to $x,y$ in different $V_{i}$. For $x,y$ in the same $\Gamma_{i}$, we have the following expansion instead. Here and below we see $\Gamma_{i}$ as a subset of $\Gamma$, and elements of $\Gamma_{i}$ have word length 0 or $1$ in $\Gamma$.
@@
Note that $p(y,y)=0$ if there are no self-loops, i.e. $p_{e}=0$, and that we do not assume $p(y,x)\neq 0$, i.e. $x,y$ may not be neighbors.
@@
Let $n_{i}=|\Gamma_{i}|$. To each $\Gamma_{i}$ we associate the $\nu_{i}=n_{i}(n_{i}-1)$ functions $z\mapsto\zeta_{x}^{z}(y)$ for $(x,y)\in\Gamma_{i}\times\Gamma_{i}$ with $x\neq y$. [^3] Actually we have $\zeta_{x}^{z}(y)=\zeta_{e}^{z}(yx^{-1})$, so there are only $n_{i}$ distinct functions, but for our argument, there is no need to use this reduction. Ordering them arbitrarily, we denote them by $\zeta_{i,j,k}$, where $i\in[m]$, $j\in[n_{i}]$ and $k\in[n_{i}-1]$. Equation (6.9) thus gives the following finite system of quadratic equations, of which $\{h_{i,j,k}=\zeta_{i,j,k}^{z}\}$ (with $i\in[m]$, $j\in[n_{i}]$, and $k\in[n_{i}-1]$) is a solution: for each $i\in[m]$, $j\in[n_{i}]$, and $k\in[n_{i}-1]$,
@@
The fact that $R^{z}(e,e)$ is algebraic appeared in a number of works, in various degrees of generality and justification [8, 43, 77], sometimes by simply exhibiting a system of algebraic equations that $R^{z}(e,e)$ satisfies. This is not enough in general, for example $h_{1}+zh_{2}=0$ and $2h_{1}+2zh_{2}=0$ has a solution $h_{1}(z)=ze^{z}$ and $h_{2}(z)=-e^{z}$ which is not algebraic.
@@
The spectrum $\sigma(\lambda(p))$ has a non-trivial absolutely continuous part. There exists a discrete set $D_{0}\subset\mathbb{R}$ such that for any $E\in\mathbb{R}\setminus D_{0}$, $i\leq m$ and $x,y\in\Gamma_{i}$, the function $\eta\mapsto\zeta_{x}^{E+i\eta}(y)$ has a finite limit as $\eta\rightarrow 0$, and the functions $E\mapsto\zeta_{x}^{E+i0}(y)$ are analytic on $\mathbb{R}\setminus D_{0}$. Furthermore, the maps $E\in\sigma(\lambda(p))\setminus D_{0}\mapsto R^{E+i0}(e,e)$ and $E\in\sigma(\lambda(p))\setminus D_{0}\mapsto\zeta_{x}^{E+i0}(y)$ have finitely many zeroes.
@@
Next, we know from Proposition 6.6 and the Newton-Puiseux theorem [69, Theorem 3.5.2] that there exists a discrete set $D_{1}\subset\mathbb{R}$ such that for any $i\leq m$ and $x,y\in\Gamma_{i}$, the map $z\mapsto\zeta_{x}^{z}(y)$ is analytic near any $z_{0}\notin D_{1}$ (the function has moreover an expansion as a Puiseux series near $z_{0}\in D_{1}$, and the set $D_{1}$ is necessarily real because we already know that $\zeta_{x}^{z}(y)$ is analytic on $\mathbb{C}\setminus\mathbb{R}$). The function $\zeta_{x}^{z}(y)$ is also nonzero on $\mathbb{C}\setminus\mathbb{R}$ by definition. It follows from analyticity that the zeroes of $\zeta_{x}^{z}(y)$ are in some discrete set $D_{1}\cup D_{2}$, thus inside the compact set $\sigma(\lambda(p))$ there are only finitely many of them.
We next prove the statement for $R^{z}$. As the function $E\mapsto\sum_{y}p(x,y)\zeta_{x}^{E+i0}(y)-E$ is analytic on $\mathbb{R}\setminus D_{1}$, its zeroes form a discrete set $D_{3}$, and its intersection with $\sigma(\lambda(p))$ which is compact is finite. By (6.4), we deduce that $E\mapsto R^{E+i0}(e,e)$ is analytic on $\mathbb{R}\setminus(D_{1}\cup D_{3})$. Note that $R^{E+i0}(e,e)$ is necessarily nonzero on $\mathbb{R}\setminus D_{1}$ by (6.4). ∎
@@
The spectrum $\sigma(\lambda(p))$ of a free product as above consists of a union of intervals $J_{j}$ of purely absolutely continuous spectrum, plus possibly finitely many eigenvalues, and (1.5) holds in any closed interval $I_{0}$ contained in the interior of some $J_{j}$.
Arguing as in [7] one can prove that the absolutely continuous part is in fact a finite union of intervals, but we shall not pursue this here.
@@
Following the notation used in the proof of Proposition 6.7, let $F=D_{1}\cup D_{2}\cup D_{3}$. Since $E\mapsto R^{E+i0}(e,e)$ is analytic on $\mathbb{R}\setminus F$, the set $M^{ac}=\{E\in\mathbb{R}\setminus F:0<\Im R^{E+i0}(e,e)<\infty\}$ is non-empty, and open. Thus, $M^{ac}$ is a union of open intervals, and the absolutely continuous part $\overline{M^{ac}}$ of the spectrum is a union of intervals $J_{j}$. The eigenvalues of $\lambda(p)$ are poles of $E\mapsto\Im R^{E+i0}(e,e)$, which can only occur if $E\in F$, and since $\lambda(p)$ is bounded, $\sigma(\lambda(p))\cap F$ is a finite set. By definition and continuity of $E\mapsto\Im R^{E+i0}(e,e)$, we know that (1.5) holds true on each closed interval $I_{0}$ contained in the interior of some $J_{j}$. Such $I_{0}$ is compact as $\sigma(\lambda(p))$ is bounded. ∎
@@
In this section, we check that (1.15) holds. We first prove an elementary lemma which shows that in order to prove the $4$-th moment bound (1.15) in regions where (1.5) holds, it is sufficient to check the vanishing of a quantity involving a supremum, instead of a sum.
@@
$\sum_{g\in\Gamma}\eta^{2}(\Im R^{\lambda_{0}+i\eta})(e,g)^{4}|g|^{C_{1}^{\prime}}\leq\Bigl(\sup_{g\in\Gamma}\eta|(\Im R^{\lambda_{0}+i\eta})(e,g)|^{2}|g|^{C_{1}^{\prime}}\Bigr)\sum_{g\in\Gamma}\eta(\Im R^{\lambda_{0}+i\eta})(e,g)^{2}.$
@@
Let us prove (6.11) for free products. We first show that there exists $\delta>0$ such that for any $E\in I$, any $\eta$ small enough, any distinct $i_{1},i_{2}\in[m]$, and any $j_{1},k_{1}\in[n_{i_{1}}],j_{2},k_{2}\in[n_{i_{2}}]$ with $j_{\ell}\neq k_{\ell}$ for $\ell=1,2$, there holds
@@
$\omega_{\underline{\ell}}^{z}=\omega_{\ell_{r}}^{z}\ldots\omega_{\ell_{1}}^{z}.$
Finally for $s=0,1,2$, we pick $g_{s}\in\Gamma_{i_{s}}$ such that $\zeta^{z}_{i_{s},j_{s},k_{s}}=\zeta^{z}_{e}(g_{s})$ (for this, we first pick $x_{s},y_{s}$ such that $\zeta^{z}_{i_{s},j_{s},k_{s}}=\zeta^{z}_{x_{s}}(y_{s})$ and then let $g_{s}=y_{s}x_{s}^{-1}$). Then
$\omega_{\underline{\ell}}^{z}=\zeta_{e}^{z}(g_{\ell_{r}})\zeta_{e}^{z}(g_{2})\zeta_{e}^{z}(g_{\ell_{r-1}})\zeta_{e}^{z}(g_{2})\cdots\zeta_{e}^{z}(g_{\ell_{1}})\zeta_{e}^{z}(g_{2}).$
Since $\ell_{i}\in\{0,1\}$, it follows that $g_{\ell_{r}}g_{2}g_{\ell_{r-1}}g_{2}\ldots g_{\ell_{1}}g_{2}\in\Gamma$, where $g_{2}$ appears every other time. Therefore we deduce by Lemma 6.4 and the Ward identity (2.16)
@@
Write any $y\in\Gamma$ as $y=g_{i_{r},j_{r}}\ldots g_{i_{1},j_{1}}$ where $g_{i_{\ell},j_{\ell}}\in\Gamma_{i_{\ell}}$ and $i_{\ell}\neq i_{\ell+1}$ for any $\ell\leq r-1$. Set $L(y)=r$. We have $L(y)\geq|y|/M$ where $M=\sup_{i}|\Gamma_{i}|$ and recall that $|y|$ denotes the word metric of $y\in\Gamma$. Combining (6.12) with Lemma 6.4, we obtain